Consider the production function y= x^(1/2). However, assume that, for political reasons, it is not feasible to hire fewer then one unit of input. Thus, assume x>= 1. (a) Show that the inverse production function x(y) is convex. (b) The price of y is p= 10. Find the firm's marginal product MP(x) and average product AP(x). (c) Find the firm's value of marginal product VMP(x) and value of average product VAP(x). (d) Find the firm's input demand curve x*(w). (e) Suppose the price of x is w=1. Calculate the firm's profit.
Answer #1

Answer.

A. Inverse production function is x=y^2.

Differentiating withrespect to y.

dx/dy = 2y

Repeating the same

dx^2/dy^2 = 2 >0

Which shows that invesei produciopr function is convex.

B. Marginal product =1/2√x.

And average product = 1/√x.

C. VMP = 5/√x. (MP*P)

And VAP = 10/√x.

D. For input function we maximize profit

p*y-wx

Differentiating with respect to x.

p/2√x-w=0

For p=10.

5/w= √x

So, x = 25/w^2.

As HOMEWORKLIB guidelkgui first four parts are answered.